### v1

@@ 1 Introduction
The current picture of the Universe and its evolution history is based on the concordance $\Lambda$ cold dark matter ($\Lambda$CDM) model, which describes gravitational phenomena according to the laws of general relativity and assumes the Universe to be statistically homogeneous and isotropic on cosmological scales. This simple model is capable of fitting a wide variety of observations: the temperature and polarisation anisotropies of the cosmic microwave background (Hinshaw et al. 2013; Planck Collaboration: Aghanim, N. et al. 2020), the clustering of galaxies on large scales (Alam et al. 2021; DESI Collaboration et al. 2024), weak gravitational lensing, which causes distortion in the observed shapes of galaxies (Abbott et al. 2022), and the distance measurements from supernovae (Scolnic et al. 2018). However, the standard model is not fully satisfactory. From a theoretical point of view, it requires that the majority of the energy content in our Universe takes some unknown form: a cosmological constant, whose inferred value lacks fundamental understanding, drives the current accelerated expansion of the Universe, while the matter content is mostly composed of a non-baryonic component that is yet to be directly detected. Furthermore, as recent cosmological experiments have become more precise, some tensions between parameters inferred from different probes have emerged (Di Valentino et al. 2021a, b; Schöneberg et al. 2022; DESI Collaboration et al. 2024; Tutusaus et al. 2023). Currently, it remains unclear whether these discrepancies either reveal a breakdown of the $\Lambda$CDM model, a sign that astrophysical or systematic effects have been overlooked in the analyses in tension, or whether they are nothing more than a statistical fluke.
@@
The Euclid Spectroscopic Survey will provide a catalogue of more than 25 million emission-line galaxies using slitless spectroscopy for redshift determination, in the redshift range between $z=0.9$ and $z=1.8$. One of the key measurements that Euclid will carry out from these data is that of the growth rate at different cosmic times, using redshift-space distortions (RSDs; Kaiser 1987). RSDs manifest because we do not directly observe galaxies where they are but rather infer their physical distances from their observed redshifts, assuming that the latter is approximately given by the Hubble expansion. However, the observed redshift is also affected by the peculiar velocities of galaxies through the Doppler shift, for example. Since we do not know the peculiar velocities of individual objects, we cannot correct each redshift individually but rather model the distortions induced by the peculiar motions in the 2-point statistics. The main effect is that galaxies moving along the line-of-sight toward an overdensity will appear closer to each other. Therefore, the observed 2-point correlation function (or its Fourier counterpart, the power spectrum) is not isotropic, but also includes even multipoles (quadrupole and hexadecapole) that can be measured to extract cosmological information (Alam et al. 2016; Sanchez et al. 2017).
@@
More recently, a fully relativistic computation has shown that in addition to the Kaiser RSDs, typically exploited in anisotropic galaxy clustering analysis, there are other subtle physical effects that can also influence the observed distribution of galaxies, such as gravitational redshift, Doppler effects, Shapiro time delay and the integrated Sachs–Wolfe effect (Yoo et al. 2009; Yoo 2010; Bonvin & Durrer 2011; Challinor & Lewis 2011; Jeong et al. 2012). Although their impact on even multipoles of the correlation function has been shown to be negligible even for future galaxy surveys (Jelic-Cizmek et al. 2021), it is possible to isolate them using multiple tracers. Importantly, while standard RSDs generate only even multipoles of the correlation function at leading order, the gravitational redshift and further Doppler-type contributions generate odd multipoles in the 2-point cross-correlation function (McDonald 2009; Bonvin et al. 2014).
@@
Measuring the odd multipoles is interesting, as they provide complementary information to the even multipoles. In particular, the dipole is sourced in part by the gravitational redshift, so in principle is sensitive to the distortion of time (Sobral-Blanco & Bonvin 2022). Measuring the distortion of time through the dipole, in combination with existing probes, would thus allow for a range of novel tests. These include: a test of the validity of the equivalence principle for dark matter, by combining with RSDs (Bonvin & Fleury 2018), a test of the validity of local position invariance (Saga et al. 2023), a novel way to measure the anisotropic stress, by combining with weak gravitational lensing (Sobral-Blanco & Bonvin 2021; Tutusaus et al. 2023) and more generally a way to distinguish between modifications of gravity and additional forces in the dark sector (Castello et al. 2022; Bonvin & Pogosian 2023; Castello et al. 2024). Furthermore, the dipole and octupole can be used to improve our understanding of astrophysical properties of galaxies (Sobral-Blanco et al. 2024).
@@ 2 Relativistic effects in galaxy clustering
The fluctuations in galaxy counts have been computed in a perturbed spacetime assuming linear perturbations in Yoo et al. (2009)Yoo (2010),  Bonvin & Durrer (2011)Challinor & Lewis (2011) and  Jeong et al. (2012), and to higher orders in perturbation theory in Yoo & Zaldarriaga (2014)Bertacca et al. (2014),  Di Dio et al. (2014a)Nielsen & Durrer (2017)Di Dio & Seljak (2019) and Di Dio & Beutler (2020). The leading order contributions relevant for this work are:
$\Delta(\boldsymbol{n},z)=b\,\delta-\frac{1}{\mathcal{H}}\partial_{r}(\boldsymbol{V}\cdot\boldsymbol{n})\\ +\frac{1}{c\,\mathcal{H}}\partial_{r}\Psi+\left(1-\frac{\mathcal{H}^{\prime}}{\mathcal{H}^{2}}+\frac{(5s-2)\,c}{r\mathcal{H}}-5s+f^{\rm evo}\right)\frac{\boldsymbol{V}}{c}\cdot\boldsymbol{n}+\frac{1}{\mathcal{H}}\frac{\boldsymbol{V}^{\prime}}{c}\cdot\boldsymbol{n}$ (3)
@@
Equation (3) does not include lensing magnification, nor subdominant relativistic effects that are proportional to the gravitational potentials – the full expression can be found in  Challinor & Lewis (2011), Bonvin & Durrer (2011) and Di Dio et al. (2013). These extra relativistic effects are suppressed compared to the gravitational redshift and Doppler corrections, and therefore can be safely neglected (Bonvin 2014). Lensing magnification also contributes to the number counts, and is expected to be an important effect for both the photometric and the spectroscopic Euclid Surveys (Euclid Collaboration: Lepori et al. 2022; Euclid Collaboration: Jelic-Cizmek et al. 2023), and the cross-correlation of photometric galaxy clustering with the cosmic microwave background (Bermejo-Climent et al. 2020). However, the impact of lensing magnification on the dipole of the 2-point correlation function becomes relevant only at high redshift and large scales, so remains negligible for most applications; see, for example, Lepori et al. (2020).
@@ 2.1 Relativistic dipole: linear regime
The information contained in the galaxy counts is normally compressed into summary statistics. For a spectroscopic survey, the most common data compression consists of estimating the multipoles of the 2-point correlation function (2PCF) or the multipoles of the power spectrum in Fourier space. In real space, where only the galaxy overdensity contributes to the galaxy counts, the correlation function is isotropic as a consequence of the cosmological principle. The second term in the first line of Eq. (3), known as Kaiser RSDs (Kaiser 1987), breaks the isotropy of the correlation function. At leading order, RSDs generate only even multipoles of the correlation function: quadrupole and hexadecapole. The relativistic effects in the second line of Eq. (3) or Eq. (5) break the symmetry of the correlation function when two tracers with different biases are cross-correlated (McDonald 2009; Bonvin et al. 2014). Therefore, by measuring the odd multipoles in the cross-correlation function, we can isolate the relativistic effects from the dominant density and Kaiser RSDs contributions. The main contribution to the antisymmetric part of the cross-correlation function is given by the dipole and is the focus of this work.
@@ 2.2 Relativistic dipole: nonlinear regime
Assuming that we cross-correlate two populations having a typical potential depth $\overline{\Psi}^{0}_{\rm B}$ for the bright population and $\overline{\Psi}_{\rm F}$ for the faint population, the full dipole signal is modelled as a sum of a linear part, given by Eq. (8), and a nonlinear contribution,
@@
Together with the model of the nonlinear potential presented in Sect. 3.2, the model in Eq. (14) has been shown to reproduce the measured dipole in $N$-body simulations down to $d\approx 5\,h^{-1}\,{\rm Mpc}$ very well (Saga et al. 2020). Since the contribution from the nonlinear gravitational redshift given in Eq. (15) is suppressed as the separation $d$ increases, the total signal in Eq. (14) recovers the results of linear theory on large scales.
@@
In Fig. 1 we show the nonlinear model for the dipole at the redshifts considered in this work. The signal is largely boosted on scales below $30\,h^{-1}\,{\rm Mpc}$ by the contribution of the nonlinear gravitational redshift.
@@ 2.3 Theory covariance
${\rm Cov}^{(\rm th)}={\rm Cov}^{(\rm noise)}+{\rm Cov}^{(\rm SV)}+{\rm Cov}^{({\rm SV}\,\times\,{\rm noise})}\,$ (18)
@@ 3.1.1 The Flagship galaxy mock
The main data product was produced on the fly during the simulation and is a continuous full-sky particle light cone that extends to $z=3$. The 3D particle light-cone data were used to identify roughly 150 billion dark matter haloes using Rockstar (Behroozi et al. 2013), and to create all-sky 2D maps of dark matter in 200 tomographic redshift shells between $z=0$ and $z=99$. The halo catalogue and the set of 2D dark matter maps are the main inputs for the Flagship mock galaxy catalogue. For a detailed description of the production of the catalogue, see Euclid Collaboration: Castander et al. (2024), but the main steps can be summarised as follows. First, galaxies were generated following a combination of halo occupation distribution (HOD) and abundance matching techniques. Following the HOD prescription, haloes were populated with central and satellite galaxies. Each halo contains a central galaxy and a number of satellites that depends on the halo mass. The halo occupation was chosen to reproduce observational constraints of galaxy clustering in the local Universe (Zehavi et al. 2011).
The luminosities of the central galaxies were assigned by performing abundance matching between the halo mass function of the simulation halo catalogue and the galaxy luminosity function (LF). The satellite luminosities were assigned assuming a universal Schechter LF in which the characteristic luminosity depends on the central luminosity in a way that ensures that the global LF agrees with observations. Galaxies were divided into three colour types, namely red, green, and blue, and the central and satellite galaxies in each group were distributed to match the clustering as a function of colour observed by Zehavi et al. (2011). The radial positions of the satellites within a given halo follow a best-fit ellipsoidal Navarro–Frenk–White (NFW) profile (Navarro et al. 1997). Finally, we assign galaxy lensing properties within the Born approximation following the ‘onion universe’ approach presented in Fosalba et al. (2008) and Fosalba et al. (2015).
@@ 3.1.2 Selection of spectroscopic sample
Table 1: Properties of the full catalogue. The objects have been selected using the redshift provided by the Flagship data, which includes Doppler effects but neglects the gravitational redshift. Fluxes are not magnified, although flux magnification has a very small impact on this estimate. The estimates of bias and local count slope are described in Sect. 3.5.
@@ 3.2 Implementation of gravitational redshift in Flagship
In FS, the observed redshift is estimated from the peculiar velocities of galaxies. Other relativistic corrections, including gravitational redshift and the integrated Sachs–Wolfe effect, are neglected. In this work, to include the gravitational redshift $z_{\Psi}=-\Psi/c^{2}$, we have simply corrected the observed redshift $z_{\mathrm{obs,FS}}$ provided in the Flagship catalogue by
@@
To estimate the gravitational redshift we require the potential $\Psi$ at each galaxy. Since on small scales the relativistic dipole has been found to be primarily sourced by gravitational redshift (Breton et al. 2019), it is important that the potential is accurately modelled on these scales. But given that high-resolution gravitational potential maps are not provided by the Flagship simulation, we will adopt a semi-analytical approach and assume that each halo where a galaxy is found can be modelled using a NFW density profile. The gravitational potential of the halo can then be estimated from the Poisson equation. The NFW gravitational potential $\Psi\equiv\Psi_{\mathrm{NFW}}$ at a comoving distance $R$ from the centre of the halo is
@@
Here, $\rho_{\rm s}$ and $R_{\rm s}$ are the scale density and the scale radius, the two parameters that describe a NFW density profile. They depend on the concentration parameter, the virial radius, and the virial overdensity of the halo, see Mo et al. (2010) and Saga et al. (2022) for the exact expression of these quantities. For satellite galaxies, at a comoving distance $R$ from the centre of the parent halo, we can estimate the gravitational potential from Eq. (25). For the central galaxies, we can take the limit

### v2

@@ 1 Introduction
The current picture of the Universe and its evolution history is based on the concordance $\Lambda$ cold dark matter ($\Lambda$CDM) model, which describes gravitational phenomena according to the laws of general relativity and assumes the Universe to be statistically homogeneous and isotropic on cosmological scales. This simple model is capable of fitting a wide variety of observations: the temperature and polarisation anisotropies of the cosmic microwave background (Hinshaw et al. 2013; Planck Collaboration: Aghanim, N. et al. 2020), the clustering of galaxies on large scales (Alam et al. 2021; DESI Collaboration et al. 2024), weak gravitational lensing, which causes distortion in the observed shapes of galaxies (Abbott et al. 2022), and the distance measurements from supernovae (Scolnic et al. 2018; DES Collaboration et al. 2024). However, the standard model is not fully satisfactory. From a theoretical point of view, it requires that the majority of the energy content in our Universe takes some unknown form: a cosmological constant, whose inferred value lacks fundamental understanding, drives the current accelerated expansion of the Universe, while the matter content is mostly composed of a non-baryonic component that is yet to be directly detected. Furthermore, as recent cosmological experiments have become more precise, some tensions between parameters inferred from different probes have emerged (Di Valentino et al. 2021a, b; Schöneberg et al. 2022; DESI Collaboration et al. 2024; Tutusaus et al. 2023). Currently, it remains unclear whether these discrepancies either reveal a breakdown of the $\Lambda$CDM model, a sign that astrophysical or systematic effects have been overlooked in the analyses in tension, or whether they are nothing more than a statistical fluke.
@@
The Euclid Spectroscopic Survey will provide a catalogue of more than 25 million emission-line galaxies using slitless spectroscopy for redshift determination, in the redshift range between $z=0.9$ and $z=1.8$. One of the key measurements that Euclid will carry out from these data is that of the growth rate at different cosmic times, using redshift-space distortions (RSDs; Kaiser 1987). RSDs manifest because we do not directly observe galaxies where they are but rather infer their physical distances from their observed redshifts, assuming that the latter is approximately given by the Hubble expansion. However, the observed redshift is also affected by the peculiar velocities of galaxies through the Doppler shift, for example. Since we do not know the peculiar velocities of individual objects, we cannot correct each redshift individually but rather model the distortions induced by the peculiar motions in the two-point statistics. The main effect is that galaxies moving along the line-of-sight toward an overdensity will appear closer to each other. Therefore, the observed two-point correlation function (or its Fourier counterpart, the power spectrum) is not isotropic, but also includes even multipoles (quadrupole and hexadecapole) that can be measured to extract cosmological information (Alam et al. 2016; Sanchez et al. 2017).
@@
More recently, a fully relativistic computation has shown that in addition to the Kaiser RSDs, typically exploited in anisotropic galaxy clustering analysis, there are other subtle physical effects that can also influence the observed distribution of galaxies, such as gravitational redshift, Doppler effects, Shapiro time delay, and the integrated Sachs–Wolfe effect (Yoo et al. 2009; Yoo 2010; Bonvin & Durrer 2011; Challinor & Lewis 2011; Jeong et al. 2012). Although their impact on even multipoles of the correlation function has been shown to be negligible even for future galaxy surveys (Jelic-Cizmek et al. 2021), it is possible to isolate them using multiple tracers. Importantly, while standard RSDs generate only even multipoles of the correlation function at leading order, the gravitational redshift and further Doppler-type contributions generate odd multipoles in the two-point cross-correlation function (McDonald 2009; Bonvin et al. 2014).
@@
Measuring the odd multipoles is interesting, as they provide complementary information to the even multipoles. In particular, the dipole is sourced in part by the gravitational redshift, so in principle is sensitive to the distortion of time (Sobral-Blanco & Bonvin 2022). Measuring the distortion of time through the dipole, in combination with existing probes, would thus allow for a range of novel tests. These include: a test of the validity of the equivalence principle for dark matter, by combining with RSDs (Bonvin & Fleury 2018), a test of the validity of local position invariance (Saga et al. 2023), a novel way to measure the anisotropic stress, by combining with weak gravitational lensing (Sobral-Blanco & Bonvin 2021; Tutusaus et al. 2023), and more generally a way to distinguish between modifications of gravity and additional forces in the dark sector (Castello et al. 2022; Bonvin & Pogosian 2023; Castello et al. 2024). Furthermore, the dipole and octupole can be used to improve our understanding of astrophysical properties of galaxies (Sobral-Blanco et al. 2024).
@@ 2 Relativistic effects in galaxy clustering
The fluctuations in galaxy counts have been computed in a perturbed spacetime assuming linear perturbations in Yoo et al. (2009)Yoo (2010),  Bonvin & Durrer (2011)Challinor & Lewis (2011), and  Jeong et al. (2012), and to higher orders in perturbation theory in Yoo & Zaldarriaga (2014)Bertacca et al. (2014),  Di Dio et al. (2014a)Nielsen & Durrer (2017)Di Dio & Seljak (2019), and Di Dio & Beutler (2020). The leading order contributions relevant for this work are:
$\Delta(\boldsymbol{n},z)=b\,\delta-\frac{1}{\mathcal{H}}\partial_{r}(\boldsymbol{V}\cdot\boldsymbol{n})\\ +\frac{1}{c\,\mathcal{H}}\partial_{r}\Psi+\left(1-\frac{\mathcal{H}^{\prime}}{\mathcal{H}^{2}}+\frac{(5s-2)\,c}{r\mathcal{H}}-5s+f^{\rm evo}\right)\frac{\boldsymbol{V}}{c}\cdot\boldsymbol{n}+\frac{1}{\mathcal{H}}\frac{\boldsymbol{V}^{\prime}}{c}\cdot\boldsymbol{n}\,.$ (3)
@@
Equation (3) does not include lensing magnification, nor subdominant relativistic effects that are proportional to the gravitational potentials – the full expression can be found in  Challinor & Lewis (2011), Bonvin & Durrer (2011), and Di Dio et al. (2013). These extra relativistic effects are suppressed compared to the gravitational redshift and Doppler corrections, and therefore can be safely neglected (Bonvin 2014). Lensing magnification also contributes to the number counts, and is expected to be an important effect for both the photometric and the spectroscopic Euclid Surveys (Euclid Collaboration: Lepori et al. 2022; Euclid Collaboration: Jelic-Cizmek et al. 2024), and the cross-correlation of photometric galaxy clustering with the cosmic microwave background (Bermejo-Climent et al. 2020). However, the lensing contribution to the dipole is negligible for most applications, because the only contribution comes from an integral over the correlation scale. Bonvin et al. (2014) show that in linear theory the impact of lensing on the dipole is largely subdominant with respect to the Doppler and gravitational redshift at $z=2$; this is confirmed to be valid in the nonlinear regime, using $N$-body simulations, in Breton et al. (2019).
@@ 2.1 Relativistic dipole: Linear regime
The information contained in the galaxy counts is normally compressed into summary statistics. For a spectroscopic survey, the most common data compression consists of estimating the multipoles of the two-point correlation function (2PCF) or the multipoles of the power spectrum in Fourier space. In real space, where only the galaxy overdensity contributes to the galaxy counts, the correlation function is isotropic as a consequence of the cosmological principle. The second term in the first line of Eq. (3), known as Kaiser RSDs (Kaiser 1987), breaks the isotropy of the correlation function. At leading order, RSDs generate only even multipoles of the correlation function: quadrupole and hexadecapole. The relativistic effects in the second line of Eq. (3) or Eq. (5) break the symmetry of the correlation function when two tracers with different biases are cross-correlated (McDonald 2009; Bonvin et al. 2014). Therefore, by measuring the odd multipoles in the cross-correlation function, we can isolate the relativistic effects from the dominant density and Kaiser RSDs contributions. The main contribution to the antisymmetric part of the cross-correlation function is given by the dipole and is the focus of this work.
@@ 2.2 Relativistic dipole: Nonlinear regime
Assuming that we cross-correlate two populations having a typical potential depth $\overline{\Psi}^{0}_{\rm B}$ for the bright population and $\overline{\Psi}^{0}_{\rm F}$ for the faint population, the full dipole signal is modelled as a sum of a linear part, given by Eq. (8), and a nonlinear contribution,
@@
Together with the model of the nonlinear potential presented in Sect. 3.2, the model in Eq. (14) has been shown to reproduce very well the measured dipole in $N$-body simulations down to $d\approx 5\,h^{-1}\,{\rm Mpc}$ (Saga et al. 2020). Since the contribution from the nonlinear gravitational redshift given in Eq. (15) is suppressed as the separation $d$ increases, the total signal in Eq. (14) recovers the results of linear theory on large scales.
@@
In Fig. 1 we show the nonlinear model for the dipole at the redshifts considered in this work. The signal is largely boosted on scales below $30\,h^{-1}\,{\rm Mpc}$ by the contribution of the nonlinear gravitational redshift. In our model, density and velocity fields are assumed to be linear. We have tested the impact of nonlinearities in the matter field by replacing the linear matter power spectrum in Eq. (16) with the nonlinear power spectrum. In all scales of interest for our work, the difference between these two cases is negligible.
Other models have been proposed in the literature, which use perturbation theory and include higher-order bias correction (Di Dio & Seljak 2019; Di Dio & Beutler 2020; Beutler & Di Dio 2020). The main disadvantage of this perturbative approach is that it introduces extra parameters into the model that cannot be directly measured in simulations but must instead be fitted to the simulated data vectors. The model presented in Saga et al. (2022) and outlined in this section is more predictive, primarily because the extra ingredients $\overline{\Psi}^{0}_{\rm B}$ and $\overline{\Psi}^{0}_{\rm F}$ can be estimated from simulations.
@@ 2.3 Theory covariance
${\rm Cov}^{(\rm th)}={\rm Cov}^{(\rm noise)}+{\rm Cov}^{(\rm SV)}+{\rm Cov}^{({\rm SV}\,\times\,{\rm noise})}\,.$ (18)
@@ 3.1.1 The Flagship galaxy mock
The main data product are produced on the fly during the simulation and is a continuous full-sky particle light cone that extends to $z=3$. The 3D particle light-cone data are used to identify roughly 150 billion dark matter haloes using Rockstar (Behroozi et al. 2013), and to create all-sky 2D maps of dark matter in 200 tomographic redshift shells between $z=0$ and $z=99$. The halo catalogue and the set of 2D dark matter maps are the main inputs for the Flagship mock galaxy catalogue. For a detailed description of the production of the catalogue, see Euclid Collaboration: Castander et al. (2024), but the main steps can be summarised as follows. First, galaxies are generated following a combination of halo occupation distribution (HOD) and abundance matching techniques. Following the HOD prescription, haloes are populated with central and satellite galaxies. Each halo contains a central galaxy and a number of satellites that depends on the halo mass. The halo occupation is chosen to reproduce observational constraints of galaxy clustering in the local Universe (Zehavi et al. 2011).
The luminosities of the central galaxies are assigned by performing abundance matching between the halo mass function of the simulation halo catalogue and the galaxy luminosity function (LF). The satellite luminosities are assigned assuming a universal Schechter LF in which the characteristic luminosity depends on the central luminosity in a way that ensures that the global LF agrees with observations. Galaxies are divided into three colour types, namely red, green, and blue, and the central and satellite galaxies in each group were distributed to match the clustering as a function of colour observed by Zehavi et al. (2011). The radial positions of the satellites within a given halo follow a best-fit ellipsoidal Navarro–Frenk–White (NFW) profile (Navarro et al. 1997). Finally, galaxy lensing properties are assigned within the Born approximation following the ‘onion universe’ approach presented in Fosalba et al. (2008) and Fosalba et al. (2015).
@@ 3.1.2 Selection of spectroscopic sample
Table 1: Properties of the full catalogue.
@@
333 The objects have been selected using the redshift provided by the Flagship data, which includes Doppler effects but neglects the gravitational redshift. Fluxes are not magnified, although flux magnification has a very small impact on this estimate. The estimates of bias and local count slope are described in Sect. 3.5.
@@ 3.2 Implementation of gravitational redshift in Flagship
In FS, the observed redshift is estimated from the peculiar velocities of galaxies. Other relativistic corrections, including gravitational redshift and the integrated Sachs–Wolfe effect, are neglected. In this work, to include the gravitational redshift $z_{\Psi}=-\Psi/c^{2}$, we simply corrected the observed redshift $z_{\mathrm{obs,FS}}$ provided in the Flagship catalogue by
@@
To estimate the gravitational redshift we require the potential $\Psi$ at each galaxy. Since on small scales the relativistic dipole is primarily sourced by gravitational redshift (Breton et al. 2019), it is important that the potential is accurately modelled on these scales. But given that high-resolution gravitational potential maps are not provided by the Flagship simulation, we adopted a semi-analytical approach and assumed that each halo where a galaxy is found can be modelled using a NFW density profile. The gravitational potential of the halo can then be estimated from the Poisson equation. The NFW gravitational potential $\Psi\equiv\Psi_{\mathrm{NFW}}$ at a comoving distance $R$ from the centre of the halo is
@@
Here, $\rho_{\rm s}$ and $R_{\rm s}$ are the scale density and the scale radius, the two parameters that describe a NFW density profile. They depend on the concentration parameter, the virial radius, and the virial overdensity of the halo, see Mo et al. (2010) and Saga et al. (2022) for the exact expression of these quantities. For satellite galaxies, at a comoving distance $R$ from the centre of the parent halo, we estimated the gravitational potential from Eq. (25). For the central galaxies, we took the limit